OpenStudy (anonymous):

A 30-year-old worker plans to retire at age 65. He believes that $500,000 is needed to retire comfortably. How much should be deposited now at 7% compounded monthly to meet the $500,000 retirement goal? (Do not round until the final answer. Then, round to the nearest cent.)

OpenStudy (rsadhvika):

Compound interest

OpenStudy (rsadhvika):

\(\huge A = P(1 +\frac{r}{n})^{n*t}\)

OpenStudy (rsadhvika):

A = 500,000 r = .07 n = 12 t = 65-30 = 35

OpenStudy (rsadhvika):

\(\huge 500,000 = P(1+\frac{.07}{12})^{12*35}\) you can solve for P ?

OpenStudy (anonymous):

Thank you :)

OpenStudy (rsadhvika):

wc !(: